you are dealt a 5 card hand. whats is the probability that the hand you have is a straight (aces are high)?
Answer by Edwin McCravy(20054)

We will find the number of straights including straight flushes
and then subtract the number of straight flushes, in case you
don't want to include those:

We choose the denominations any of these 10 ways:
A,2,3,4,5
2,3,4,5,6
3,4,5,6,7
4,5,6,7,8
5,6,7,8,9
6,7,8,9,10
7,8,9,10,J
8,9,10,J,Q
9,10,J,Q,K
10,J,Q,K,A

For each of those ways we choose the suit for
the lowest card in 4 ways. That's 10*4 ways
For each of those ways we choose the suit for
the next to lowest card in 4 ways. That's 10*4*4
For each of those ways we choose the suit for
the middle card in 4 ways. That's 10*4*4*4
For each of those ways we choose the suit for
the next to highest card in 4 ways. That's 10*4*4*4*4
For each of those ways we choose the suit for
the highest card in 4 ways. That's 10*4*4*4*4*4
ways or 10*4^5 = 10240

We now find the number of straight flushes 
(including Royal flushes).

We choose the denominations any of these 10 ways:
A,2,3,4,5
2,3,4,5,6
3,4,5,6,7
4,5,6,7,8
5,6,7,8,9
6,7,8,9,10
7,8,9,10,J
8,9,10,J,Q
9,10,J,Q,K
10,J,Q,K,A
We can choose the 1 suit they must all have to be 
a straight flush in 4 ways.

That's 10*4 or 40 straight flushes.

So the number of straights is 10240-40 = 10200

The number of 5-card poker hands is
52 cards choose 5 = 52C5 = 2598960.

So the probability of being dealt a straight that
is not a straight flush is 

10200/2598960 = 5/1274 (dividing top and bottom by 2040).

That's about 0.003924646781789638932496075353218210361067503924646781789

If you want the probability including the straight
flushes, then it's

10240/2598960 = 128/32487 (dividing top and bottom by 80).

That's about 0.003940037553482931634192138393819066087973650998861082894…

[They are the same rounded to 4 decimal places. 0.0039.  That's 
because straight flushes are so rare].

Edwin